OpenStudy (andijo76):

one card is selected from a deck of playing cards. Determine the probability of selecting a card greater than 9 or a black card.

OpenStudy (anonymous):

The deck has 52 cards. Cards greater than 9 is all 10's, Jacks, Queens, Kings and Aces. That sums up to 20 cards. 4 of these are black. Half of the deck is black. That's 26 cards. So the probability is 20+16=36 out of 52 cards.

OpenStudy (anonymous):

Wrong numbers. The probability would be 16+26/52=>42/52

OpenStudy (andijo76):

thanks

Directrix (directrix):

Of the red cards in the deck, 10 have value greater than 9 (5 diamonds and 5 hearts). There are 26 black cards in the deck. P(selecting a card greater than 9 or a black card) = 10/52 + 26/52 = 36/ 52 = 9/13.